OpenStudy (anonymous):

According to Boyles LAw the pressure "P" of a compressed gas is inversely proportional to the volume, "V". IF a pressure of 20 pounds per square inch exist when the volume of the gas is 500 cubic inches what exist when it is 400 cubic incvhes

OpenStudy (anonymous):

P=kV so 20=k*500 k=1/25 so 1/25*400=16 answer is 16

OpenStudy (anonymous):

and it is pounds per square inch